Suppose that in a certain region of space there is an electrostatic field and also a magnetostatic field. Show that, although the Poynting vector can be nonzero, the surface integral of media%2F0f6%2F0f68f8f8-86f1-4408-9135-f4is annulled on a closed surface within this region, which does not contain charges.
Answer #1

The Poynting vector can be written as

\vec S=\vec E\times \vec B

where I have neglected the 1/\mu_o term for brevity as that is a constant term and I will not need it for the proof

The closed surface integral can be written as follows using the divergence theorem

\oint_A (\vec S\cdot \hat n) dA=\int_V (\nabla \cdot \vec S)dV

where the surface integral is over closed surface A and the volume integral is over the volume enclosed in the surface A. The divergence of S can be bewritten as

\nabla \cdot \vec S=\nabla \cdot(\vec E\times \vec B)=\vec B\cdor (\nabla \times \vec E)-\vec E(\nabla \times \vec B)

Now note that for static field configuration and the space without free charges, using the Maxwell's eqations

\nabla \times \vec E=-\frac{\partial \vec B}{\partial t}=0

because E and B are static fields(means they are independent of time)

And

\nabla \times \vec B=\frac{1}{c^2\epsilon_o}\vec J+\frac{1}{c^2}\frac{\partial \vec E}{\partial t}=0

Because there is no free current in the enclosed space over which the integral is performed and E is static

Therefore

\nabla \cdot \vec S=0

and hence the closed surface integral of the Poynting vector within a region, which does not contain charges or currents is

\oint_A (\vec S\cdot \hat n) dA=\int_V (\nabla \cdot \vec S)dV=0